1. Given the demand function Q=15-3P, what is the quantity demanded if the price of the good is PHP 10.00



2. Given that


P↓0=2, P↓1=3


Q↓0=15, Q↓1=30


Solve the price elasticity of demand coefficien. What type of elasticity is it?

1
Expert's answer
2021-11-04T13:40:25-0400

1. Q = 15 - 3×10 = -15 units, so at this price zero quantity will be demanded.

2. The price elasticity of demand coefficient is:

"Ed = \frac{30 - 15} {2-3} ×\frac{2+3} {30+15} = -5/3 = -1.67."

So, the demand is elastic.